# Robust Optimization-based Autonomous Dynamic Soaring with a Fixed-Wing UAV

### v1

@@ IV-C Path Following Control
Once a path has been generated, the vehicle tracks it using a dynamic-inversion based controller. Dynamic inversion requires analytic invertibility and differential flatness of the model. However, for a fixed-wing gliding aircraft, tracking of any arbitrary trajectory in state space cannot be achieved due to a lack of degrees of freedom to generate thrust and direct yaw control. Therefore, we employ a projection of the flat outputs onto a feasible subspace, which can be tracked by the aircraft. This projection allows us to make use of differential flatness based control methodologies for path following without encountering infeasible configurations.
@@ IV-C1 Position Dynamics
We construct projected flatness relations for tracking a position reference $\xi=\bm{\mathbf{r}}_{\text{ref}}(t)$. The required force ${\bm{\mathbf{f}}}_{\text{ref}}$ to fulfill $\ddot{\bm{\mathbf{r}}}=\ddot{\bm{\mathbf{r}}}_{\text{ref}}$ is, in general, obtained by simply plugging $\ddot{\bm{\mathbf{r}}}_{\text{ref}}$ into the Newton-Euler equation
@@
For dynamic inversion, we must equate (2) with (12), which, however, is infeasible due to the coupling between lift and drag. Instead, we only consider the perpendicular component to the path, which is necessary to track lateral kinematics. We decompose the required net force into its lift and drag components, i.e.
@@ IV-C2 Attitude Dynamics
To compensate for modeling errors and disturbances, the sensor-based control approach incremental nonlinear dynamic inversion is chosen to control both the position and attitude of the aircraft. We choose to employ a cascaded control approach inspired by [Tal] for position and attitude control of the gliding aircraft. The cascaded control architecture is shown in Fig. 6.
@@ IV-C3 Position Control
In the position controller, the position, velocity, and acceleration reference and related errors are used to compute an attitude reference for the attitude controller. First, a linear PD control law is applied to the position and velocity errors to obtain an acceleration command for the actual system. Similar to [Tal], the control law can be written as
@@ IV-C4 Attitude Control
In the attitude controller, the reference values for the attitude $\bm{\mathbf{q}}_{\text{ref}}$ and its temporal derivatives ${\bm{\omega}}_{\text{ref}}$ and $\dot{\bm{\omega}}_{\text{ref}}$ and the related errors are used to compute control surface deflections for the system. The structure of the attitude controller is identical to the structure shown in [Tal]. However, the dynamic inversion is performed differently due to the zero sideslip assumption. In a first step, the error angle vector $\bm{\zeta}$ is obtained as
$\begin{split}\bm{\mathbf{q}}_{\text{e}}&=\bm{\mathbf{q}}^{-1}\circ\bm{\mathbf{q}}_{\text{c}}\\ \bm{\zeta}&=\frac{2\arccos({q}_{\text{e}}^{w})}{1-({q}_{\text{e}}^{w})^{2}}\begin{bmatrix}{q}_{\text{e}}^{x}\quad{q}_{\text{e}}^{y}\quad{q}_{\text{e}}^{z}\end{bmatrix}^{T}\text{,}\end{split}$ (24)
where $\bm{\mathbf{q}}_{\text{e}}=\begin{bmatrix}{q}_{\text{e}}^{w}\quad{q}_{\text{e}}^{x}\quad{q}_{\text{e}}^{y}\quad{q}_{\text{e}}^{z}\end{bmatrix}^{T}$ is the attitude error quaternion and $\bm{\mathbf{q}}$ is the current attitude quaternion. The feed-forward terms ${\bm{\omega}}_{\text{ref}}$ and $\dot{\bm{\omega}}_{\text{ref}}$ are derived from the on-path (closest point) target attitude required to follow the path with the current inertial speed and wind. The error angle vector $\bm{\mathbf{\zeta}}$ is then used in a linear PD control law, to obtain an angular acceleration command for the system. The linear control law is written as
@@
where $\mathbf{K}_{\text{P}}^{\text{att}}$ and $\mathbf{K}_{\text{P}}^{\text{att}}$ are diagonal control gain matrices corresponding to the proportional and derivative terms, respectively. The diagonal entries of these matrices are selected according to the control authority available around the respective body axis. Analogously to the position controller, in a second step, the moment increment $\Delta\mathbf{M}$ is computed and added to the current modeled moment to create a moment command $\mathbf{M}_{\text{c}}$
@@
where $\mathbf{M}_{\text{model}}$ is the current moment computed using the model in Eq. (5). Also, it is assumed that the angular term is slowly varying and is identical to its low-pass filtered counterpart [Tal]. To compute $\mathbf{M}_{\text{model}}$, the current actuator deflection needs to be known. Howver, due to a lack of position feedback, we approximate the servo as first order system and estimate the deflection based on a corresponding low-pass filter applied to the actuator commands. Finally, the moment command is used to compute actuator deflection references for ailerons and elevator using the flatness relation (21).
Note that the rudder deflection is allocated independently to maintain zero side-slip. This is achieved through a turn coordination. Based on the current speed and lateral acceleration readings from the IMU, it computes the required body-$z$ angular rate $r^{*}$ to maintain zero side-slip. The ideal yaw rate $r^{*}$ is then passed through an airspeed scaling and fed to the rudder, together with a linear feedback term
@@ V-A Validation Approach
Validating a dynamic soaring framework across a broad range of wind-conditions is practically challenging, due to limited flight testing time and often hard-to-predict wind conditions for planning effective field trips. With this work, we thus focus on evaluating the components of the framework individually, but stress important aspects for successful, integrated real-world dynamic soaring. First, (robust) performance under a controllable range of wind-conditions and estimation errors is best assessed in simulation. Second, to ensure that the simulation results transfer to reality, we need to validate the wind-estimator and demonstrate that the system can i) closely follow paths as considered in the simulation and ii) that it exhibits a similar energy evolution, i.e. one as modeled and optimized for. This step-by-step evaluation is presented in the following.
V-B Simulation
@@
An example of the resulting responses can be seen in Fig. 9. The simulation results indicate that the proposed architecture can achieve robust dynamic soaring under unknown, static deviations in the wind field.
@@
V-C Flight Experiments
Hardware experiments were conducted with a powered (folding-prop) model glider aircraft with a wingspan of $1.8\text{\,}\mathrm{m}$ and a mass of $1.35\text{\,}\mathrm{k}\mathrm{g}$. Wind estimation and path following control were integrated into the PX4 Autopilot, running on a Pixhawk 4 flight management unit. The pre-computed trajectory primitives were stored on the SD card.
TABLE I: Parameter values used in the experiments.
Param $m$ $A_{\text{wing}}$ $C_{L0}$ $C_{L1}$ $C_{D0}$ $C_{D1}$ $C_{D2}$
Value 1.35 0.4 0 0 0.0169 -0.0940 0.3580
V-C1 Wind Field Estimation
The wind estimation module from Fig. 7 (local and global estimator) was evaluated in a flight test in light shear conditions. While it is difficult to obtain accurate ground-truth measurements of the wind field, we focus on the overall plausibility and consistency of the estimated wind field. For this purpose, the unmanned aerial vehicle was manually piloted in a shear layer behind a mountain ridge, altering altitude to cover the vertical distribution. The estimates show variation in the local measurements across heights (Fig. 8), which results from the coarse local estimation model (6) and the spatial and temporal variations in the wind field. Regardless, the global parameter estimates converge to a near-static equilibrium.
(a) Path following arbitrary ellipse path
(b) Path following strong winds
Figure 10: Path following results with (a) arbitrary inclined ellipse paths and electric propulsion (b) strong winds in Chasseral.
V-C2 Robust Path Following
In the first flight experiment, the capability of the control architecture to track inclined paths with high-acceleration turns was evaluated. To imitate paths tracked in a dynamic soaring scenario, a set of inclined ellipses of different sizes and inclination angles were used as a reference path. In addition, the path of one robust soaring trajectory obtained from (11) (for an arbitrary wind field and initial velocity) was added to the set of paths. The test was conducted using the onboard electric motor set at a constant throttle to produce sufficient thrust to maintain flight in low-wind conditions ($\textbf{w}||\leq 3\frac{m}{s}$), while switching between the reference paths. The resulting flight path and tracking error is shown in Fig. 10. While the covered state trajectories in this experiment do not represent dynamic soaring, the provided reference paths demand aggressive maneuvering, similar to dynamic soaring. The control architecture demonstrates good path following, with smooth switching between reference paths. The resulting tracking error largely remained below $3\text{\,}\mathrm{m}$, apart from the initial convergence phase after switching references.
To demonstrate the tracking performance of the controller subject to strong gusting, a similar experiment was also carried out in strong wind behind a ridge (Fig. 1, Fig. 10). The present wind field offered wind speeds between $8\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ and $12\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, with gusts up to $14\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$. The results highlight the ability of the control architecture to deliver solid tracking performance at airspeeds up to $32\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
The performance of the tracking controller is crucial because precise maneuvering at high speed is needed for satisfactory performance in unknown and rapidly changing wind conditions (as when crossing the shear layer). It should be noted that the controller performance is independent of an accurate global wind estimate since only the instantaneous, local wind estimates are used for wind correction, while the global estimates are only used for planning purposes.
The control performance further demonstrates small sim-to-real discrepancy with respect to path following precision.
V-C3 Closed Loop Efficiency
@@
Finally, we need to verify that close tracking of a desired path, as demonstrated in the previous experiments, also results in the expected energy evolution on the real system. This is to verify that the sim-to-real gap is also small with respect to potential losses from, e.g., unknown drag terms or aggressive maneuvering. To this end, we compare the airspeed evolution in unpowered flight, where the system tracks an aggressive spiral-like trajectory. The flight is performed in low-wind conditions to focus on model-error induced effects, noting that robustness to wind estimation errors is shown in simulation. The airspeed evolution is compared to the simulated case in Fig. 11. Note that the airspeed is sensitive to the aerodynamic performance during this test, as illustrated by two simulated trajectories with perturbed drag coefficients that differ significantly from the nominal case with the expected parameters. Thus, any model-error or control-induced loss of airspeed is expected to become visible during the test. Results, however, show a close match between the real flight test and the nominal simulation case. This indicates that the closed-loop performance of the pipeline closely matches the simulation. Together with the simulation-based demonstration of robustness against wind-estimation errors, this completes the element-wise validation of the proposed pipeline.
@@ VI CONCLUSIONS
In this article, we propose and demonstrate a framework for robust autonomous dynamic soaring. We show that, by using a optimization approach which considers wind uncertainty, the dynamic soaring approach can be more robust against wind uncertainty. The optimized trajectories are precomputed, and selected based on the estimated wind conditions. We demonstrate an incremental nonlinear dynamic inversion-based path following controller that can reliably track these paths. We believe that this is a step towards demonstrating autonomous dynamic soaring on an unmanned aerial vehicle.

### v2

@@ IV-C Path Following Control
Once a path is generated, the vehicle tracks it using a dynamic-inversion based controller. However, for a gliding aircraft, full differential flatness is unattainable due to the lack of thrust and direct yaw authority. We therefore project the flat outputs onto a feasible subspace, allowing a path following approach without encountering infeasible configurations.
@@ IV-C1 Position Dynamics
We construct projected flatness relations for tracking a position reference $\xi=\bm{\mathbf{r}}_{\text{ref}}(t)$. The required force ${\bm{\mathbf{f}}}_{\text{ref}}$ to fulfill $\ddot{\bm{\mathbf{r}}}=\ddot{\bm{\mathbf{r}}}_{\text{ref}}$ is obtained from the Newton-Euler equation
@@
Direct inversion via (2) is infeasible due to lift-drag coupling. Instead, we project the requirement onto the path-perpendicular vector to satisfy lateral kinematics, decomposing the net force as
@@ IV-C2 Attitude Dynamics
To compensate for modeling errors and disturbances, the sensor-based control approach incremental nonlinear dynamic inversion is chosen to control both the position and attitude of the aircraft. We choose to employ a cascaded control approach inspired by [21] for position and attitude control of the gliding aircraft. The cascaded control architecture is shown in Fig. 7.
@@ IV-C3 Position Control
In the position controller, the position, velocity, and acceleration reference and related errors are used to compute an attitude reference for the attitude controller. First, a linear PD control law is applied to the position and velocity errors to obtain an acceleration command for the actual system. Similar to [21], the control law can be written as
@@ IV-C4 Attitude Control
In the attitude controller, the reference values for the attitude $\bm{\mathbf{q}}_{\text{ref}}$ and its temporal derivatives ${\bm{\omega}}_{\text{ref}}$ and $\dot{\bm{\omega}}_{\text{ref}}$ and the related errors are used to compute control surface deflections for the system. The structure of the attitude controller is identical to the structure shown in [21]. However, the dynamic inversion is performed differently due to the zero sideslip assumption. In a first step, the error angle vector $\bm{\zeta}$ is obtained as
$\bm{\zeta}=\frac{2\arccos({q}_{\text{e}}^{w})}{1-({q}_{\text{e}}^{w})^{2}}\begin{bmatrix}{q}_{\text{e}}^{x}\quad{q}_{\text{e}}^{y}\quad{q}_{\text{e}}^{z}\end{bmatrix}^{T}\text{,}\quad\bm{\mathbf{q}}_{\text{e}}=\bm{\mathbf{q}}^{-1}\circ\bm{\mathbf{q}}_{\text{c}}$ (23)
where $\bm{\mathbf{q}}_{\text{e}}=\begin{bmatrix}{q}_{\text{e}}^{w}\qquad{q}_{\text{e}}^{x}\qquad{q}_{\text{e}}^{y}\qquad{q}_{\text{e}}^{z}\end{bmatrix}^{T}$ is the attitude error quaternion and $\bm{\mathbf{q}}$ is the current attitude quaternion. The feed-forward terms ${\bm{\omega}}_{\text{ref}}$ and $\dot{\bm{\omega}}_{\text{ref}}$ are derived from the on-path (closest point) target attitude required to follow the path with the current inertial speed and wind. The error angle vector $\bm{\mathbf{\zeta}}$ is then used in a linear PD control law, to obtain an angular acceleration command for the system. The linear control law is written as
@@
where $\mathbf{K}_{\text{P}}^{\text{att}}$ and $\mathbf{K}_{\text{P}}^{\text{att}}$ are diagonal control gain matrices corresponding to the proportional and derivative terms, respectively. Analogously to the position controller, the moment increment $\Delta\mathbf{M}$ is computed and added to the current modeled moment to create a moment command $\mathbf{M}_{\text{c}}$
@@
where $\mathbf{M}_{\text{model}}$ is the current moment computed using the model in Eq. (5). Also, it is assumed that the angular term is slowly varying and identical to its low-pass filtered counterpart [21]. To obtain an estimate of $\mathbf{M}_{\text{model}}$, we approximate the servo as first order system and estimate the deflection based on a corresponding low-pass filter applied to the actuator commands. Finally, the moment command is used to compute actuator deflection references for ailerons and elevator using the flatness relation (20).
Note that the rudder deflection is allocated independently to enforce zero side-slip. This is achieved through turn coordination, where the required body-$z$ angular rate $r^{*}$ to maintain zero side-slip is computed based on the current speed and lateral acceleration. The desired yaw rate $r^{*}$ is then passed to the rudder, together with a linear feedback term
@@ V EXPERIMENTAL RESULTS
Validating the proposed approach across a broad range of wind-conditions is challenging, due to the stochastic nature of wind. Therefore, we focus on evaluating components of the framework individually, but stress important aspects for successful real-world dynamic soaring. First, performance under a controllable range of wind-conditions and estimation errors is best assessed in simulation. Second, to ensure that the simulation results transfer to reality, we demonstrate that the system can i) closely follow paths as considered in the simulation and ii) that it exhibits a similar energy evolution one has modeled and optimized for.
V-A Simulation
We simulate the closed-loop behavior of the trajectory optimizer and the path-tracking controller under wind estimation errors to evaluate the robustness of the pipeline. We consider a nominal wind field ($\bar{w}_{max}=$ $12\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, $\bar{s}=$ $0.3\text{\,}\mathrm{m}$, $\bar{h}_{ref}=$ $0\text{\,}\mathrm{m}$) which is used by the trajectory optimizer alongside 4 perturbed wind fields ($M=5$) with a $25\%$ relative perturbation of the parameters $\bar{w}_{max}$ and $\bar{s}$. The actual wind field parameters that are used in simulation are sampled from a uniform distribution with a $25\%$ relative perturbation radius of $\bar{w}_{max}$ and $\bar{s}$. We use the Dryden model to add turbulence the mean wind field. Dryden model parameters $L$ and $\sigma$ were derived by fitting the model to the residual error signal from subtracting a Gaussian Process-modeled mean wind field from the flight data in Fig. 10 (b). To further account for the worst-case turbulence, we increase the standard deviation to over-approximate 99.99%, obtaining $L_{u}=$36\text{\,}\mathrm{m}$\$, $L_{v}=$20\text{\,}\mathrm{m}$\$, $L_{w}=$10\text{\,}\mathrm{m}$\$, $\sigma_{u}=$0.96\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$, $\sigma_{v}=$0.84\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$, $\sigma_{w}=$0.57\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$\$. We further add a time-varying, random estimation bias of $1\text{\,}\mathrm{m}$ to the local wind estimate. During each rollout, the perturbed wind-field parameters are held constant and overlaid with the turbulence, while the global wind field for trajectory selection is kept at the nominal values. Trajectory switching due to changes in initial speed takes place at the end of a path segment. Note that, while the reference path being tracked has been optimized based on the nominal wind-field, the tracking controller has access to the actual, sampled wind field. This reflects the proposed use of a global and local wind-estimator. Robustness was evaluated over 8-cycle simulations, with success being completion without airspeed dropping below the $10\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ stall/tracking limit. Despite wind perturbations and modeling uncertainties, the architecture achieved a 97.3% success rate and shows continuous increase of air speed in Fig. 9.
Figure 9: Simulation of dynamic soaring in perturbed wind fields with overlayed turbulence from a Dryden model. The UAV follows robust soaring paths optimized for a nominal wind field with ($|\mathbf{w}_{0}$||= $12\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, $s=$ $0.3\text{\,}\mathrm{m}$, $h_{ref}=$ $0\text{\,}\mathrm{m}$) and 4 perturbed wind fields with $25\%$ relative perturbations on $\mathbf{w}_{0}||$ and $s$ each. In the simulation, the wind field parameters are perturbed by random additive, uniform noise with $25\%$ relative perturbation radius on $\mathbf{w}_{0}||$ and $s$. The simulation is performed for 300300 runs (only 1010 shown for visual clarity).
V-B Flight Experiments
Hardware experiments were conducted with a model glider Multiplex EasyGlider with a wingspan of $1.8\text{\,}\mathrm{m}$ and a mass of $1.35\text{\,}\mathrm{k}\mathrm{g}$. Wind estimation and path following control were integrated into the PX4 Autopilot, running on a Pixhawk 4 flight management unit. The pre-computed trajectory primitives were stored on the SD card of the flight management unit, and the air speed was measured using a pitot tube. The position and attitude controllers operate at $50\text{\,}\mathrm{Hz}$ and $200\text{\,}\mathrm{Hz}$, respectively.
TABLE II: Parameter values used in the experiments.
Param $m$ $A_{\text{wing}}$ $C_{L0}$ $C_{L1}$ $C_{D0}$ $C_{D1}$ $C_{D2}$ $C_{A0}$ $C_{A1}$ $C_{B1}$
Value 1.35 0.4 0.356 2.354 0.029 0.378 1.984 0.279 4.313 -3.464
V-B1 Wind Field Estimation
The wind estimation module was evaluated in a flight test in light shear conditions. While it is difficult to obtain accurate ground-truth measurements of the wind field, we focus on the overall plausibility and consistency of the estimated wind field. For this purpose, the unmanned aerial vehicle was manually piloted in a shear layer behind a mountain ridge, altering altitude to cover the vertical distribution. The estimates show variation in the local measurements across different heights (Fig. 8), which results from the coarse local estimation model (6) and spatial and temporal variations in the wind field. Regardless, the global parameter estimates converge to a near-static equilibrium.
(a) Path following arbitrary ellipse path (Zürich, Switzerland)
(b) Path following strong winds (Chasseral, Switzerland)
Figure 10: Vehicle trajectory colored with tracking error with ground track as grey (a) arbitrary inclined ellipse paths and electric propulsion (b) strong winds. Vehicle attitude is enlarged 5 times for visibility.
V-B2 Robust Path Following
@@
V-B3 Closed Loop Efficiency
@@
Finally, we verify that close tracking of a desired path also results in the expected energy evolution on the real system. This verifies that the sim-to-real gap is also small with respect to potential losses from, e.g., unknown drag terms or aggressive maneuvering. We compare the airspeed evolution in unpowered flight, where the system tracks an aggressive spiral-like trajectory. The flight is performed in low-wind conditions to focus on model-error induced effects, noting that robustness to wind estimation errors is shown in simulation. The airspeed evolution is compared to the simulated case in Fig. 11. Note that the airspeed is sensitive to the aerodynamic performance during this test, as illustrated by two simulated trajectories with perturbed drag coefficients that differ significantly from the nominal case with the expected parameters. Thus, any model-error or control-induced loss of airspeed is expected to become visible during the test. Results, however, show a close match between the real flight test and the nominal simulation case. This indicates that the closed-loop performance of the pipeline closely matches the simulation.
@@ VI CONCLUSIONS
In this article, we propose and demonstrate a framework for robust autonomous dynamic soaring. We show that, by using a optimization approach which considers wind uncertainty, the dynamic soaring approach can be more robust against wind uncertainty. The optimized trajectories are precomputed, and selected based on the estimated wind conditions. We demonstrate an incremental nonlinear dynamic inversion-based path following controller that can reliably track these paths. We believe that this is a step towards demonstrating autonomous dynamic soaring on an unmanned aerial vehicle. A complete in-the-loop hardware validation of the proposed framework are left for future work due to the need for an optimized experimental platform. Further, the integration of secondary objectives like distance coverage and topographic constraints are relevant points for future work.